# Ultra-Short Pulse Biphoton Source in Lithium Niobate Nanophotonics at 2 µm

### v1

@@
Abstract
@@ 1 Introduction
Fig. 1: A diagram of the source operation. Femtosecond pump pulses are injected into a dispersion-engineered periodically polled lithium niobate (PPLN) waveguide to produce an ultra-fast photon pair via type-0 spontaneous parametric down-conversion (SPDC). A) An illistration of the spontaneous parametric down converison (SPDC) process used to generate photon pairs. B) 2-photon microscopy image of the polled region and an SEM of a representative device. C) The theoretical and measured SPDC spectra.
@@
Since the first observations of the quantum nature of light, considerable efforts have gone into building systems gaining a quantum advantage over purely classical technologies to benefit a variety of fields including sensing, metrology, communication, and computation[1, 2, 3]. While many physical platforms have been studied to exploit quantum phenomenon, photonics stands out as it offers several benefits including room temperature operation and scalability. Integrated photonics [4, 5] allows the realization of many devices in a chipscale footprint, paving the way for the development of large-scale integrated systems for QIP at room temperature [6, 7].
@@
In recent years, thin-film lithium niobate (TFLN) has emerged as a favorable platform for the development of entangled photon sources given its high $\chi^{(2)}$. Additionally, dispersion engineering enables the generation of broadband entangled photon pairs with short coherence times. Broadband operation is important in a variety of quantum applications as it leads to improved sensitivity in optical coherence tomography [8], greater spectral coverage for quantum spectroscopy [9, 10], sharp temporal behavior for atomic state control [11], and lithographic resolution exceeding the classical diffraction limit [12]. Wavelength-multiplexing schemes [13, 14] as well as high-dimensional entanglement protocols from quantum networks [15, 13] can also make good use of broadband sources.
@@
There have been many successful demonstrations of broadband biphoton sources in TFLN [16, 17, 18]. However, these devices are still limited to the telecom wavelength range. Moreover, they have been demonstrated in the continuous-wave (CW) regime and have been unsuitable for ultrafast operation. Operation in the ultrafast regime requires near-zero group velocity dispersion (GVD) conditions for both the pump and signal wavelengths, as well as near-zero group velocity mismatch (GVM) between the pump and signal. Ultrashort-pulse sources of quantum states of light are particularly important because of opportunities in time-division multiplexing and the enhancement of nonlinear interactions as a result of increased peak power. By time-division multiplexing with well-defined nanosecond time bins, large-scale quantum states are realized using fiber-based systems [19, 20]. Their extension to the ultrafast femtosecond time bins can enable generation and manipulation of similarly large-scale quantum states in integrated photonics [21]. Stronger nonlinear interactions as a result of temporal confinement are also an important step towards all-optical non-Gaussian sates and operations [22, 23]. While our implementation currently relies on a table-top mode-locked laser, significant progress has been made in integrating short pulse sources on chip, especially in LN nanophotonics, with a 4.8-ps 10-GHz chip-scale mode-locked laser demonstrated in [24] and a chip-scale electro-optic comb source with a 520-fs duration and a 30-GHz repetition rate demonstrated in [25].
@@
Dispersion engineering in TFLN using the waveguide geometry provides new opportunities for classical and quantum ultrafast photonics beyond broadband CW operation. Particularly, achieving near-zero group-velocity mismatch between the signal and pump along with zero group-velocity dispersion at the signal and pump wavelengths simultaneously has enabled record-breaking gain-bandwidth products [26], ultralow-energy ultrafast all-optical switching [27] and the on-chip generation and measurement of broadband squeezed states[28]. Unlike the previous nanophotonic biphoton sources [29, 30, 31, 16, 32, 33, 34, 35, 36, 37], we utilize a near-zero dispersion regime for realization of ultrashort-pulse photon pairs in nanophotonics. Operating in the 2-µm band in lithium niobate is particularly advantageous as it yields better fabrication tolerances compared with similar geometries for the 1550-nm band [38], and low GVD for both pump and signal simultaneously as well as the ability to match their group velocities and create strong temporal confinement for generated photon pairs [39].
@@
Biphoton sources beyond the telecommunication band are also important for quantum key distribution (QKD) applications. For instance, recent work on free-space QKD at 1550 nm has made a compelling case for moving to longer wavelengths to avoid limitations caused by solar irradiance [40], which decreases at longer wavelengths [41]. Integrated photonics platforms can also benefit from operation at longer wavelengths because of lower scattering losses[42] and more efficient high-speed modulators[43, 44]. Classical and quantum networks are becoming practical in the 2-µm window given the progress on sources [45, 46], as well as low-dispersion and low-nonlinearity photonic crystal fibers [47, 48], and high-speed detectors [49]. Along with thulium/holmium amplifiers offering 40 THz of bandwidth [50], these technologies can facilitate the construction of wide area classical and quantum networks in the mid-IR. Quantum effects and secure communications have also been successfully demonstrated in the 2-µm band with Hong-Ou-Mandel visibility of 88.1% measured in bulk PPLN by [33] and polarization-based quantum key distribution implemented by [51] with a key rate of 0.254 bits/pair.
@@
In this work, we present the first ultrafast biphoton source in nanophotonics PPLN capable of supporting ultra-short pulse propagation at both the pump and signal/idler wavelengths (Fig.1). Our source covers 17 THz of 3-dB bandwidth centered at 2.09 µm, and produces a 165-fs biphoton wavepacket at the output. Through 2 and 3-fold coincidence counting experiments, we demonstrate a coincidence-to-accidentals ratio (CAR) of 945, a pair generation rate of 8.8 GHz/mW or 440 kHz/mW/GHz, and a heralded $g_{H}^{(2)}(0)$ of 0.027, all of which are state-of-the-art for this wavelength range in nanophotonics. Combined with the recent advances in ultrafast lithium niobate nanophotonics [26, 28, 27, 25], and the wide variety of high-performance components [5], our source demonstrates a practical path towards ultrafast on-chip QIP.
@@
Fig. 2: A) The quasi-TE waveguide mode for pump light at 1.045 µm. B) The quasi-TE waveguide mode for signal light at 2.09 µm. The waveguide top width is 2.5 µm. C) Dispersion profile for the pump wavelength. D) Dispersion profile for the signal wavelength. E) Group velocity mismatch at different signal wavelengths relative to 1.045 µm.
@@
Fig. 3: A) The experimental setup used to characterize the source. MLL is a 250-MHz 75-fs mode-locked laser centered at 1.045 µm. HWP is a half-wave plate. PBS is a polarizing beam splitter. VND is a variable neutral density filter. Obj is a reflective objective. FC is a freespace to fiber coupler. LP is a low-pass filter. BP is a band-pass filter. SNSPDs are superconducting nanowire single photon detectors. 50/50 are balanced fiber beamsplitters. H, S, and I denote the heralding, signal, and idler channels respectively. B) Measurement setup for performing the Hanbury Brown-Twiss (HBT) experiment. C) Measurement setup for determining $g_{H}^{(2)}(0)$. The SEM image inset is of a representative device [26].
@@ 2 Device Design and Fabrication
Our source is designed for type-0 degenerate spontaneous parametric down-conversion (SPDC) while also optimizing the dispersion parameters for short-pulse propagation. To enjoy the benefits of short pulse operation, the group-velocity dispersion at both the pump and signal wavelengths must be minimized to maintain the temporal confinement of the pulses as they propagate along the length of the waveguide. Our waveguide design also minimizes the group-velocity mismatch between the pump and signal/idler wavelengths to minimize the temporal interaction window of the pump and generated photon pairs [52], allowing for the creation of an ultrashort signal-idler wavepacket. This short wavepacket allows for strong nonlinear interactions given the peak power enhancement as well as the use of time bins shorter than 1 ps for clock speeds exceeding 1 THz.
We target our waveguide design for conversion from a pump centered at 1.045 µm to a degenerate signal/idler pair at 2.09 µm. The dispersion curves for our design are presented in Fig.2. We achieve a GVD of 135 $fs^{2}/mm$ and 60 $fs^{2}/mm$ for the pump and signal/idler respectively, and a GVM of 33 $fs/mm$. For a 5 mm device, this GVM corresponds to a maximum temporal window for the signal/idler pair of 165 fs. This near-zero GVM and GVD regime generates a broad spectrum of signal/idler pairs as show in Fig.1 with a 3-dB bandwidth of 17 THz. In addition to temporal confinement, the waveguide geometry spatially confines both the pump and signal/idler modes to the fundamental quasi-TE mode, providing a large mode overlap to increase pair generation rate. The mode profiles are plotted in Fig.2. The 165-fs time window for our pair source is estimated based on the temporal length of the laser input pulse and the dispersion of the waveguide, which have been measured via atomic force microscopy and confirmed with mode-solver simulations. Shorter gain windows, the gain being responsible for pair generation, for similar devices have been demonstrated in [53]. The temporal length of the photon pair can be directly measured using techniques developed in [54] and [55].
@@
While not the focus of our work, many recent works on photon-pair sources focus on using dispersion engineering to minimize the number of modes present in the output pairs [56, 57, 58, 59, 60]. In the context of SPDC, minimization of the mode number is equivalent to removing the spectral correlations from the photon pairs. This creates indistinguishable photons, which are especially important for applications where multiple independent biphoton sources are combined [61, 62, 63, 64] as this indistinguishably leads to a stronger interference of single photons from different sources. Single-mode operation for the degenerate type-0 case without the use of filtering has remained elusive due to geometric constraints stemming from the shape of the phase matching function as well as an inherent inability to satisfy the group velocity criteria used in [65, 56]. Based on the geometry of the waveguide, we estimate a mode number of 16 from the singular value decomposition of the joint spectral intensity function. For an ideal experiment with perfectly achromatic optics and detectors, this mode number should not limit the performance in CAR, generation rate, or heralded $g_{H}^{(2)}(0)$ of our source. In practice, the SNSPDs used for single photon detection experience a sharp fall-off in efficiency for increasing wavelength as does the transmissivity of the SMF-28 fibers leading up to them. We therefore perform all SNSPD measurements through a 48-nm bandpass filter centered at 2.09 µm to limit the impact of achromatic detection on our measurements. This reduces our mode number to 3. Using this filter during measurement prevents us from directly observing the total brightness of the source. Therefore, using a technique similar to [29], we calculate the total brightness by computing the ratio of the measured SPDC spectrum and the filter pass band. Multiplying the measured brightness by this ratio gives the total brightness.
@@
We fabricate our devices using commercially available thin-film lithium niobate on silica wafers from NANOLN, with a 700-nm thick X-cut thin-film on 4.7-µm thick SiO2. We start by using electron beam lithography to pattern poling electrodes deposited via metal evaporation. After poling, we lithographically define and etch waveguides with dry Ar+ plasma. The inset of Fig.3 shows a 2-photon microscopy image of the poled region before etching as well as an SEM image of the waveguide facet at the edge of the chip.

### v2

# article.bib

@@
[…]
article.bib
Ultra-Short Pulse Biphoton Source in Lithium Niobate Nanophotonics at 2 µm
James Williams Department of Electrical Engineering, California Institute of Technology, Pasadena, California 91125, USA Rajveer Nehra Department of Electrical Engineering, California Institute of Technology, Pasadena, California 91125, USA Department of Electrical and Computer Engineering, University of Massachusetts Amherst, Amherst, Massachusetts, 01003, USA Department of Physics, University of Massachusetts Amherst, Amherst, Massachusetts, 01003, USA Elina Sendonaris Department of Applied Physics, California Institute of Technology, Pasadena, California 91125, USA Luis Ledezma Department of Electrical Engineering, California Institute of Technology, Pasadena, California 91125, USA Robert M. Gray Department of Electrical Engineering, California Institute of Technology, Pasadena, California 91125, USA Ryoto Sekine Department of Electrical Engineering, California Institute of Technology, Pasadena, California 91125, USA Alireza Marandi Department of Electrical Engineering, California Institute of Technology, Pasadena, California 91125, USA
1 Abstract
@@ 2 Introduction
Figure 1: A diagram of the source operation. Femtosecond pump pulses are injected into a dispersion-engineered periodically polled lithium niobate (PPLN) waveguide to produce an ultra-fast photon pair via type-0 spontaneous parametric down-conversion (SPDC). A) An illistration of the spontaneous parametric down converison (SPDC) process used to generate photon pairs. B) 2-photon microscopy image of the polled region and an SEM of a representative device. C) The theoretical and measured SPDC spectra.
@@
Since the first observations of the quantum nature of light, considerable efforts have gone into building systems for gaining a quantum advantage over purely classical technologies to benefit a variety of fields including sensing, metrology, communication, and computation[degen2017quantum, gisin2007quantum, horowitz2019quantum]. While many physical platforms have been studied to exploit quantum phenomenon, photonics stands out as it offers several benefits including room temperature operation and scalability. Integrated photonics [siew2021review, zhu2021integrated] allows the realization of many devices in a chipscale footprint, paving the way for the development of large-scale integrated systems for QIP at room temperature [wang2020integrated, pelucchi2022potential].
@@
In recent years, thin-film lithium niobate (TFLN) has emerged as a favorable platform for the development of entangled photon sources given its high $\chi^{(2)}$. Additionally, dispersion engineering enables the generation of broadband entangled photon pairs with short coherence times. Broadband operation is important in a variety of quantum applications as it leads to improved sensitivity in optical coherence tomography [abouraddy2002quantum], greater spectral coverage for quantum spectroscopy [hickmansinglephoton, whittaker2017absorption], sharp temporal behavior for atomic state control [dayan2004two], and lithographic resolution exceeding the classical diffraction limit [pavel2014quantum]. Wavelength-multiplexing schemes [wengerowsky2018entanglement, pseiner2021experimental] as well as high-dimensional entanglement protocols from quantum networks [wei2022towards, wengerowsky2018entanglement] can also make good use of broadband sources.
@@
There have been many successful demonstrations of broadband biphoton sources in TFLN [javid2021ultrabroadband, zhang2023scalable, xue2021ultrabright]. However, these devices are still limited to the telecom wavelength range. Moreover, they have been demonstrated in the continuous-wave (CW) regime and have been unsuitable for ultrafast operation given their dispersive properties. Operation in the ultrafast regime requires near-zero group velocity dispersion (GVD) conditions for both the pump and signal wavelengths, as well as near-zero group velocity mismatch (GVM) between the pump and signal. Ultrashort-pulse sources of quantum states of light are particularly important because of opportunities in time-division multiplexing and the enhancement of nonlinear interactions as a result of increased peak power. By time-division multiplexing with well-defined nanosecond time bins, large-scale quantum states are realized using fiber-based systems [asavanant2019generation, larsen2019deterministic]. Their extension to the ultrafast femtosecond time bins can enable the generation and manipulation of similarly large-scale quantum states in integrated photonics [gray202340]. Stronger nonlinear interactions as a result of temporal confinement are also an important step towards all-optical non-Gaussian sates and operations [yanagimoto2023quantum, onodera2022nonlinear]. While our implementation currently relies on a table-top mode-locked laser, significant progress has been made in integrating short pulse sources on chip, especially in LN nanophotonics, with a 4.8-ps 10-GHz chip-scale mode-locked laser demonstrated in [guo2023ultrafast] and a chip-scale electro-optic comb source with a 520-fs duration and a 30-GHz repetition rate demonstrated in [yu2022integrated].
@@
Dispersion engineering in TFLN using the waveguide geometry provides new opportunities for classical and quantum ultrafast photonics beyond broadband CW operation. Particularly, achieving near-zero group-velocity mismatch between the signal and pump along with zero group-velocity dispersion at the signal and pump wavelengths simultaneously has enabled record-breaking gain-bandwidth products [ledezma2022intense], ultralow-energy ultrafast all-optical switching [guo2022femtojoule] and the on-chip generation and measurement of broadband squeezed states [nehra2022few]. Unlike the previous nanophotonic biphoton sources [steiner2021ultrabright, zhao2020high, elkus2019generation, javid2021ultrabroadband, guo2017parametric, prabhakar2020two, signorini2021silicon, sanna2022integrated, kumar2021mid, rosenfeld2020mid], we utilize a near-zero dispersion regime for the realization of ultrashort-pulse photon pairs in nanophotonics. Operating in the 2-µm band in lithium niobate is particularly advantageous as it yields better fabrication tolerances compared with similar geometries for the 1550-nm band [kuo2022towards], and low GVD for both pump and signal simultaneously as well as the ability to match their group velocities and create strong temporal confinement for generated photon pairs [jankowski2021dispersion].
@@
Biphoton sources beyond the telecommunication band are also important for quantum key distribution (QKD) applications. For instance, recent work on free-space QKD at 1550 nm has made a compelling case for moving to longer wavelengths to avoid limitations caused by solar irradiance [liao2017long], which decreases at longer wavelengths [solartables]. Integrated photonics platforms can also benefit from operation at longer wavelengths because of lower scattering losses [hagan2016mechanisms] and more efficient high-speed modulators [cao2018high, sadiq201640]. Classical and quantum networks are becoming practical in the 2-µm window given the progress on sources [phelan2018high, yang2019high], as well as low-dispersion and low-nonlinearity photonic crystal fibers [agrell2016roadmap, richardson2016new], and high-speed detectors [ye2015alingaas]. Along with thulium/holmium amplifiers offering 40 THz of gain bandwidth [li2013thulium], these technologies can facilitate the construction of wide area classical and quantum networks in the mid-IR. Quantum effects and secure communications have also been successfully demonstrated in the 2-µm band with Hong-Ou-Mandel visibility of 88.1% measured in bulk PPLN by [prabhakar2020two] and polarization-based quantum key distribution implemented by [dada2021near] with a key rate of 0.254 bits/pair.
@@
In this work, we present the first ultrafast biphoton source in nanophotonic PPLN capable of supporting ultra-short pulse propagation at both the pump and signal/idler wavelengths (Fig.1). Our source covers 17 THz of 3-dB bandwidth centered at 2.09 µm, and produces a 165-fs biphoton wavepacket at the output. Through 2 and 3-fold coincidence counting experiments, we demonstrate a coincidence-to-accidentals ratio (CAR) of 945, a pair generation rate of 8.8 GHz/mW or 440 kHz/mW/GHz, and a heralded $g_{H}^{(2)}(0)$ of 0.027, all of which are state-of-the-art for this wavelength range in nanophotonics. Combined with the recent advances in ultrafast lithium niobate nanophotonics [ledezma2022intense, nehra2022few, guo2022femtojoule, yu2022integrated], and the wide variety of high-performance components [zhu2021integrated], our source demonstrates a practical path towards ultrafast on-chip QIP.
@@
Figure 2: A) The quasi-TE waveguide mode for pump light at 1.045 µm. Arrows at the top right denote the ordinary and extraordinary material axes. B) The quasi-TE waveguide mode for signal light at 2.09 µm. C) Dispersion profile for the pump wavelength. D) Dispersion profile for the signal wavelength. E) Group velocity mismatch at different signal wavelengths relative to 1.045 µm.
@@
Figure 3: A) The experimental setup used to characterize the source. MLL is a 250-MHz 75-fs mode-locked laser centered at 1.045 µm. HWP is a half-wave plate. PBS is a polarizing beam splitter. VND is a variable neutral density filter. Obj is a reflective objective. FC is a freespace to fiber coupler. LP is a low-pass filter. BP is a band-pass filter. SNSPDs are superconducting nanowire single photon detectors. 50/50 are balanced fiber beamsplitters. H, S, and I denote the heralding, signal, and idler channels respectively. B) Measurement setup for performing the Hanbury Brown-Twiss (HBT) experiment. C) Measurement setup for determining $g_{H}^{(2)}(0)$. The SEM image inset is of a representative device [ledezma2022intense].
@@ 3 Device Design and Fabrication
Our source is designed for type-0 degenerate spontaneous parametric down-conversion (SPDC) while also optimizing the dispersion parameters for short-pulse propagation. To enjoy the benefits of short pulse operation, the group-velocity dispersion at both the pump and signal wavelengths must be minimized to maintain the temporal confinement of the pulses as they propagate along the length of the waveguide. Our waveguide design also minimizes the group-velocity mismatch between the pump and signal/idler wavelengths to minimize the temporal interaction window of the pump and generated photon pairs [vanselow2019ultra], allowing for the creation of an ultrashort signal/idler wavepacket. This short wavepacket can facilitate strong nonlinear interactions given the peak power enhancement and enable the use of time bins shorter than 1 ps for clock speeds exceeding 1 THz. We target our waveguide design for conversion from a pump centered at 1.045 µm to a degenerate signal/idler pair at 2.09 µm. The dispersion curves for our design are presented in Fig.2. We achieve a GVD of 135 $fs^{2}/mm$ and 60 $fs^{2}/mm$ for the pump and signal/idler respectively, and a GVM of 33 $fs/mm$. For comparison, bulk lithium niobate has has a GVD of 246 $fs^{2}/mm$ for pump light, a GVD of -56 $fs^{2}/mm$ for signal/idler photons, and a GVM of 115 $fs/mm$.
Our near-zero GVM and GVD regime of operation leads to a broad spectrum of signal/idler pairs as shown in Fig.1 with a 3-dB bandwidth of 17 THz. For a 5-mm long device, we achieve a maximum temporal length for the signal/idler wavepacket of 165 fs. The duration of this wavepacket can be estimated from the temporal overlap of the pump pulse with the vacuum modes into which the waveguide phase-matching and dispersion permit SPDC. As the pump propagates down the waveguide and walks off in time from the 2 µm vacuum field, an large number of time-delayed vacuum modes within the walk-off window can experience the creation of a photon pair. The total wavepacket duration is approximated from the pump pulse duration and the total walk-off time resulting from the GVM using Eq.27(a) from [wasilewski2006pulsed]. Shorter gain windows, the gain being responsible for pair generation, for similar devices have been demonstrated in [li2022all]. The temporal length of the photon pair can be directly measured using techniques developed in [sensarn2010generation] and [chekhova2018broadband]. Substantial GVD at either the pump or signal/idler frequencies will also temporally broaden generated photon pairs either by increasing the effective temporal gain window or by dispersing photon pairs after their creation [roeder2023measurement]. For 700-nm thick lithium-niobate on insulator (LNOI), achieving a large enough GVD necessary for the GVD to be the primary contribution to the temporal length of the photon pairs is generally difficult without resorting to extremely narrow waveguide geometries or deep etch depths. Therefore, our focus is on minimizing the GVM as this is the dominant contribution from the waveguide geometry to our temporal length. There exist other geometries for LNOI that also experience low GVM and GVD thanks to the relaxed fabrication tolerances and flatter dispersion curves around 1 and 2 µm [jankowski2021dispersion, ledezma2022intense].
We calculate the waveguide dispersion by first measuring the width and etch depth via atomic force microscopy, and then simulating the exact geometry with a mode-solver to find the effective refractive index, group velocity, and second-order dispersion. In addition to temporal confinement, the waveguide geometry spatially confines both the pump and signal/idler modes to the fundamental quasi-TE mode, providing a large mode overlap to increase the pair generation rate. The mode profiles are plotted in Fig.2A-B.
@@
While not the focus of our work, many recent works on photon-pair sources focus on using dispersion or poling domain engineering to minimize the number of temporal field-orthogonal modes [brecht2015photon] present in the output pairs [xin2022spectrally, dixon2013spectral, zhong2020quantum, mosley2008heralded, graffitti2018independent]. The structure and occupancy of these modes can be found by performing a Bloch-Messiah decomposition on the joint-spectral intensity (JSI) of the signal and idler photons [houde2023waveguided]. The JSI itself is estimated by the product of the energy conservation of the pump and the phase matching of the waveguide as a function of the signal and idler frequencies. A more detailed discussion of this process is presented in the supplementary. In the context of SPDC, minimization of the mode number is equivalent to removing the spectral correlations from the photon pairs. This creates indistinguishable photons, which are especially important for applications where multiple independent biphoton sources are combined [kok2007linear, humphreys2013linear, briegel1998quantum, spring2013boson] as this indistinguishably leads to a stronger interference of single photons from different sources. Single-mode operation for the degenerate type-0 case without the use of filtering has remained elusive due to constraints stemming from the shape of the phase matching function as well as an inherent inability to satisfy the group velocity criteria used in [grice2001eliminating, xin2022spectrally]. We estimate a mode number of 16 from the decomposition of the JSI. For an ideal experiment with perfectly achromatic optics and detectors, this mode number will not limit the performance in CAR, generation rate, or heralded $g_{H}^{(2)}(0)$ of our source. In practice, the SNSPDs used for single photon detection experience a sharp fall-off in efficiency for increasing wavelength as does the transmissivity of the SMF-28 fibers leading up to them. We therefore perform all SNSPD measurements through a 48-nm bandpass filter centered at 2.09 µm to limit the impact of achromatic detection on our measurements. This reduces our mode number to 3. Using this filter during measurement prevents us from directly observing the total brightness of the source. Therefore, using a technique similar to [steiner2021ultrabright], we calculate the total brightness by computing the ratio of the measured SPDC spectrum and the filter pass band. Multiplying the measured brightness by this ratio gives the total brightness.
@@
We fabricate our devices using commercially available thin-film lithium niobate on silica wafers from NANOLN, with a 700-nm thick X-cut thin-film on 4.7-µm thick SiO2. We start by using electron beam lithography to pattern poling electrodes deposited via metal evaporation. After poling, we lithographically define and etch waveguides with dry Ar+ plasma. The inset of Fig.1B shows a 2-photon microscopy image of the poled region before etching as well as an SEM image of the waveguide facet at the edge of the chip.